### v1

@@ Abstract
3D hand poses are an under-explored modality for action recognition. Poses are compact yet informative and can greatly benefit applications with limited compute budgets. However, poses alone offer an incomplete understanding of actions, as they cannot fully capture objects and environments with which humans interact. To efficiently model hand-object interactions, we propose HandFormer, a novel multimodal transformer. HandFormer combines 3D hand poses at a high temporal resolution for fine-grained motion modeling with sparsely sampled RGB frames for encoding scene semantics. Observing the unique characteristics of hand poses, we temporally factorize hand modeling and represent each joint by its short-term trajectories. This factorized pose representation combined with sparse RGB samples is remarkably efficient and achieves high accuracy. Unimodal HandFormer with only hand poses outperforms existing skeleton-based methods at 5$\times$ fewer FLOPs. With RGB, we achieve new state-of-the-art performance on Assembly101 and H2O with significant improvements in egocentric action recognition.
@@ 1 Introduction
The popularity of AR/VR headsets has driven interest in recognizing hand-object interactions, particularly through egocentric [14, 25] and multi-view cameras [33, 51]. Such interactions are inherently fine-grained; recognizing them requires distinguishing subtle motions and object state changes. State-of-the-art methods for hand action recognition [23, 47, 44, 67] primarily rely on multi- or single-view RGB streams, which are computationally heavy and unsuitable for resource-constrained scenarios like AR/VR.
Motivated by advancements in lightweight hand pose estimation methodologies leveraging monochrome cameras [26, 27, 45], and the evolution of low-dimensional sensor technologies [41, 35] such as accelerometers, MMG, EMG, demonstrating real-time hand pose estimation, we advocate for the utilization of 3D hand poses as an input modality for recognizing hand-object interactions. Hand poses are a compact yet informative representation that captures the motions and nuances of hand movements.
@@
Figure 1: We sample 3D hand poses at a high temporal resolution (dense) for understanding fine-grained hand motion and sparsely sample RGB frames to capture the scene semantics. 3D hand poses can be acquired from low-resolution monochrome cameras or low-dimensional sensors such as accelerometers, MMG, EEG, etc., facilitating an efficient understanding of hand-object interactions. Video frames and hand poses in the figure are from Assembly101 [51].
@@
Existing works on 3D pose-based action recognition have focused primarily on full-body skeletons [69, 43, 20, 66]. Hand poses differ fundamentally from full-body skeletons. In full-body recognition datasets [53, 40], the actions are predominantly static from a global perspective. The relative changes in joint or limb positions signify the action category, e.g. ‘hand waving.’ Conversely, hand joints typically move together for many actions and lack a static joint as a global reference [51] e.g. ‘put down toy.’ Full-body skeleton methods also benefit from modeling long-range spatiotemporal dependencies between joints [43], while this is less important for the hands as shown in Fig. 2 and in Sec. 3.
@@
However, the 3D pose alone is insufficient to encode the action for hands. Unlike the full-body case, where actions are self-contained by the sequence of poses, the hands are often manipulating objects [51, 33, 45]. Hand pose is excellent for identifying motions (verbs) but struggles with associated objects [51]. Therefore, supplementing pose data with visual context from images or videos is crucial for full semantic understanding. However, as we noted earlier, using dense RGB frames contradicts our objective and motivation for using hand poses.
@@
This work introduces HandFormer, a novel and lightweight multimodal transformer that leverages dense 3D hand-poses complemented with sparsely sampled RGB frames. To this end, we conceptualize an action as a sequence of short segments, which we refer to as micro-actions. Each micro-action comprises a dense sequence of pose frames and a single RGB frame. As every hand joint moves in close spatial proximity to each other, we opt to encode the pose sequence from a Lagrangian view [49] and track each joint as an individual entity. The pose trajectory and a single RGB frame are combined into micro-action features, which are then temporally aggregated with a transformer to predict the action label.
@@
Our contributions are: (i) We analyze the differences between hand pose and full-body skeleton actions and design a novel pose sequence encoding that reflects hand-specific properties. (ii) We propose HandFormer that takes a sequence of dense 3D hand pose and sparse RGB frames as input in the form of micro-actions. (iii) HandFormer achieves state-of-the-art action recognition performance on H2O [33] and Assembly101 [51] while unimodal HandFormer with only hand poses outperforms existing skeleton-based methods for verb recognition in Assembly101 incurring at least 5×5$\times$5 × fewer FLOPs. (iv) We experimentally demonstrate how hand poses are especially crucial for multi-view and egocentric action recognition.
@@ 2 Related Work
Video Action recognition. Video-based action recognition systems are well-developed with sophisticated 3D-CNN [58, 9, 68, 23] or Video Transformer [2, 3, 47, 42] architectures. However, they all bear significant computational expense for both feature extraction and motion handling, either explicitly in the form of optical flow [56, 9, 32] or implicitly through the architecture [2, 47]. Such designs are well-suited for high-facility research and certain cloud-based applications but are not suitable for integration into lightweight systems. To this end, efficiency in video understanding has been an active topic of research where the main focus has been model efficiency i.e., reducing expensive 3D operations [39, 68, 59, 22], quadratic complexity of attention [31, 47, 42] and dropping tokens [21, 4]. However, due to the high cost of using densly-sampled RGB frames, they can only offer limited temporal resolution, hindering fine-grained action understanding. As such, we propose to complement 3D hand poses with only sparsely-sampled RGB frames for developing a lightweight video understanding system.
@@
Skeleton-based Action Recognition. Skeleton-based action recognition has been approached with hand-crafted features [63, 61], sequence models like RNNs and LSTMs [19, 70], and CNN-based methods that either employ temporal convolution on the pose sequence [57] or transform the skeleton data into pseudo-images to be processed with 2D or 3D convolutional networks [30, 7, 20]. Recent advancements primarily come from GNN-based methods that utilize the skeletal data’s graph structure for constructing spatio-temporal graphs and performing graph convolutions [69, 71]. These methods often model functional links between joints that go beyond skeletal connectivity [38, 55] or aim to increase the spatiotemporal receptive field [43]. Self-attention and transformer-based methods have also been proposed in [48, 72, 66]. Most existing methods are tailored for datasets that involve full-body poses, while our work is dedicated to 3D hand pose data, focusing on the unique motion characteristics of hand skeletons. Although there are some methods that work with hand pose, such as [29, 36, 50], they are primarily designed for gesture recognition, which does not require explicit temporal modeling.
@@
Figure 2: Comparing skeletal changes in full-body actions from NTU RGB+D 120 [40] (left) and hand actions from Assembly101 [52] (right). Two pose frames at interval $T$ are shown. $J_{j,t}$ indicates the 3D coordinate of joint $j$ at timestep $t$, and $<J_{x},J_{y}>$ is the correlation between two such joints. Modeling the correlation between spatio-temporally distant joints can be informative in full-body pose but is unable to provide a useful action cue in hand pose.
@@
Fusing RGB with Skeleton. Pose or skeletal data can be used for cropping image patches for body parts [12], weighting RGB patches around the regions of interest [5, 6], or pooling deep CNN features [8, 1]. Projecting into common embedding space is done in [16], which enables pose distillation in [15]. Multi-stream architectures are also designed to employ separate paths for RGB and skeleton data with lateral connections between the streams [37, 64]. RGBPoseConv3D [20] is a two-stream architecture that proposes to use 3D CNNs for both RGB and pose. Similar to skeleton-based action recognition, these multi-modal approaches combining skeleton and RGB data primarily focus on full-body poses. Some approaches simultaneously perform both hand pose estimation and action recognition, using pose data to supervise the training process [65, 13]. However, it is important to note that the task of pose estimation is extremely low-level when compared to action recognition, which deals with high-level semantics. Actions can often be inferred even when the estimated poses are not highly precise [45].
3 Full-body Skeleton vs. Hand
@@
Existing skeleton-based action recognition datasets primarily consist of full-body poses where actions feature significant changes in limb positions relative to other body parts over time, strongly correlating with the action category. Capturing these changes can be achieved through long-range spatiotemporal modeling using graph convolutions with large receptive fields [43] or using self-attention [66]. Conversely, hand-pose actions show diverse movement patterns, as hands can move in arbitrary directions, often without significant changes in articulation. We also analyze pose sequences from NTU RGB+D 120 [40] and Assembly101 [51], calculating the distance covered by each joint and identifying the least and most active joints. We observe that, for full-body poses, these two representative joints show a significant difference in the distance covered, while for hand poses, they only depict a subtle distinction. The Pearson correlation coefficient is 0.93 for hand poses (indicating highly coupled joints) and 0.33 for full-body poses. Please see our Supplementary for the details of these computations and comparisons.
@@
An example showing the difference between body and hand pose is provided in Fig. 2, where we depict two frames separated by a time interval of $T$ frames. For the full body skeleton, the head joint $J_{H}$ remains static, while the right hand joint $J_{F}$ moves upward. Conversely, in the hand pose, all the joints move, including the wrist joint $J_{W}$ and index fingertip $J_{I}$. The spatiotemporal cross-correlation between the head joint in the first frame and the hand joint in the last frame ($<J_{H,0},J_{F,T}>$) is an important action cue, capturing the relative structural change. On the contrary, modeling a similar correlation between the wrist and a fingertip for hand skeleton ($<J_{W,0},J_{I,T}>$) does not provide a stronger cue than the wrist movement itself. To avoid such redundant spatiotemporally distant correlations, we propose dividing the hand pose sequence into micro-action blocks of fixed temporal length. This formulation allows for encoding short-term movements while enabling parameter sharing.
@@
Moreover, full-skeletal motion is a dominant feature in hand pose sequences, differentiating it from whole body human poses. While the location and orientation of the human body can be trivial for most action classes, this is not the case for the hands. The hands do not conform to any particular 6D pose, and more significantly, they frequently and unpredictably alter their 6D pose over time. A comprehensive analysis of this phenomenon is provided in our Supplementary. In this regard, we explicitly consider the global 6D poses of the hands during micro-action-based pose encoding.
@@
Figure 3: Overall architecture of HandFormer. An action segment is divided into $K$ micro-actions $\{M_{1},M_{2}\ldots M_{K}\}$. Each micro-action comprises a dense sequence of pose frames and a single RGB frame. The frame encoder $F$ and the trajectory encoder $T$ encode the single RGB frame and the dense poses, respectively, after which it is passed through a Multimodal Tokenizer. The modality-mixed PoseRGB tokens are then fed into a Temporal Transformer. [CLS] token represents the learnable action class token. Dotted modules are optional and only required when RGB modality is used.
@@ 4 HandFormer
Fig. 3 illustrates our proposed HandFormer, which consists of a sequence of micro-action blocks, a novel trajectory encoder, and a temporal aggregation module. The design of our HandFormer allows us to easily incorporate semantic context by sampling single RGB frames from certain micro-actions.
@@ 4.1 Micro-actions
Given an action segment containing $\mathcal{T}$ frames sampled at a certain fps, the input to our model comprises — i) a dense sequence of 3D hand poses, represented as $\mathbf{S}=\{P_{1},P_{2},\ldots,P_{\mathcal{T}}\}$, where $P_{t}=\{P_{t}^{left},P_{t}^{right}\}$ $\in\mathbb{R}^{2\times J\times 3}$ signifying the 3D coordinates of $J$ keypoints in the left and right hands, respectively, and ii) a sparse set of RGB frames sampled at intervals $\Delta f$, denoted as $\mathbf{V}=\{I_{1},I_{1+\Delta f},\ldots,I_{\mathcal{T}_{r}}\}$, where $I_{t}\in\mathbb{R}^{H\times W\times 3}$ are framewise RGB features and $\mathcal{T}_{r}=\left\lfloor\frac{\mathcal{T}-1}{\Delta f}\right\rfloor\times\Delta f$.
@@ 4.2 Trajectory Encoder
To encode dense hand-pose sequences within micro-actions, we devise a trajectory-based pose encoder, illustrated in Fig. 4. Each joint is represented by its trajectory of dimension $3\times N$, encapsulating the sequence of 3D coordinates across the $N$ pose frames of a micro-action. This yields $2\times J$ feature vectors for the $J$ joints of two hands. Each joint’s trajectory is passed through a TCN [34], whose parameters are shared for all joints. This produces $2\times J$ Local Trajectory Tokens. Additionally, the full-skeletal motion of the hand during the entire action is used as a reference through an additional token named Global Wrist Token. This token is generated by a separate TCN operating on the sequence of 6D poses of hands, indicating wrist location and hand orientation. Subsequently, a self-attention layer is applied to these trajectory tokens, preserving the temporal dimension for subsequent stages. This iterative process culminates in spatiotemporal average pooling, summarizing the hand motion of the micro-action.
@@ 4.3 Multimodal Tokenizer
In this section, we discuss incorporating sparsely sampled RGB frames into our HandFormer to better capture the scene semantics. A single frame is sampled from each micro-action, followed by generating an extended crop (1.25×1.25$\times$1.25 ×) around the hands. While the full image provides the overall scene context, the crop focuses specifically on hand-object interaction regions [10]. Features for both are separately generated using a pre-existing image encoder and then aggregated to enrich the hand-object interaction feature with scene context. The hand-object ROI crop can be obtained using an off-the-shelf HOI detector [54] or by using the corresponding hand pose projections.
@@
Our multimodal tokenizer receives the frame feature and trajectory encoding of a micro-action as input Fig. 3, and performs multi-modal feature interaction to enhance their features for better RGB and pose tokens. This involves concatenating the frame features and the pose trajectories, projecting them into a shared PoseRGB feature space using an MLP, and then splitting the PoseRGB feature into two parts, which are added to the original RGB and Pose features.

### v2

@@ Abstract
3D hand pose is an underexplored modality for action recognition. Poses are compact yet informative and can greatly benefit applications with limited compute budgets. However, poses alone offer an incomplete understanding of actions, as they cannot fully capture objects and environments with which humans interact. We propose HandFormer, a novel multimodal transformer, to efficiently model hand-object interactions. HandFormer combines 3D hand poses at a high temporal resolution for fine-grained motion modeling with sparsely sampled RGB frames for encoding scene semantics. Observing the unique characteristics of hand poses, we temporally factorize hand modeling and represent each joint by its short-term trajectories. This factorized pose representation combined with sparse RGB samples is remarkably efficient and highly accurate. Unimodal HandFormer with only hand poses outperforms existing skeleton-based methods at 5$\times$ fewer FLOPs. With RGB, we achieve new state-of-the-art performance on Assembly101 and H2O with significant improvements in egocentric action recognition.
@@ 1 Introduction
The popularity of AR/VR headsets has driven interest in recognizing hand-object interactions, particularly through egocentric [14, 26] and multiview cameras [34, 52]. Such interactions are inherently fine-grained; recognizing them requires distinguishing subtle motions and object state changes. State-of-the-art methods for hand action recognition [23, 48, 45, 67] primarily rely on multi- or single-view RGB streams, which are computationally heavy and unsuitable for resource-constrained scenarios like AR/VR. Motivated by advancements in lightweight hand pose estimation methods leveraging monochrome cameras [27, 28, 46], and the evolution of low-dimensional sensors [42, 36] such as accelerometers, MMG, EMG, demonstrating real-time hand pose estimation, we advocate the use of 3D hand poses as an input modality for recognizing hand-object interactions. Hand poses are a compact yet informative representation that captures the motions and nuances of hand movements.
@@
Figure 1: We densely sample 3D hand poses to understand fine-grained hand motions and sparsely sample RGB frames to capture the scene semantics. Our 3D hand poses are acquired from low-resolution monochrome cameras but can also come from wearable sensors, facilitating an efficient understanding of hand-object interactions. Video frames and hand poses in the figure are from Assembly101 [52].
@@
Existing works on 3D pose-based action recognition have focused primarily on full-body skeletons [69, 44, 20, 66]. Hand poses differ fundamentally from full-body skeletons. In full-body recognition datasets [53, 41], the actions are predominantly static from a global perspective. The relative changes in joint or limb positions signify the action category, e.g., ‘sitting down’. Conversely, hand joints typically move together for many actions and lack a static joint as a global reference [52], e.g., ‘put down toy’. Full-body skeleton methods also benefit from modeling long-range spatiotemporal dependencies between joints [44], while this is less important for the hands as shown in Fig. 2 and in Sec. 3.
@@
However, the 3D pose alone cannot encode the hands’ actions. Unlike the full-body case, where actions are self-contained by the sequence of poses, the hands are often manipulating objects [52, 34, 46]. Hand pose is excellent for identifying motions (verbs) but struggles with associated objects [52]. Therefore, supplementing pose data with visual context from images or videos is crucial for full semantic understanding. However, as noted earlier, using dense RGB frames contradicts our motivation for using hand poses.
@@
This paper introduces HandFormer, a novel and lightweight multimodal transformer that leverages dense 3D hand poses complemented with sparsely sampled RGB frames. To this end, we conceptualize an action as a sequence of short segments, termed micro-actions, which are analogous to words that form sentence-level complex actions. Each micro-action comprises a dense sequence of pose frames and a single RGB frame. As every hand joint moves in close spatial proximity, we encode the pose sequence from a Lagrangian view [50] and track each joint as an individual entity. HandFormer effectively models each micro-action whose features are then temporally aggregated for action classification. This formulation benefits from our observations on the spatiotemporal dynamics of hand movement and delivers strong performance. Additionally, limiting long-range spatiotemporal dependencies among joints and only using sparse RGB frame features improves computational efficiency, making it well-suited for always-on, low-compute AR/VR applications.
@@
Our contributions are: (i) We analyze the differences between hand pose and full-body skeleton actions and design a novel pose sequence encoding that reflects hand-specific properties. (ii) We propose HandFormer that takes a sequence of dense 3D hand pose and sparse RGB frames as input in the form of micro-actions. (iii) HandFormer achieves state-of-the-art action recognition on H2O [34] and Assembly101 [52]. Unimodal HandFormer with only hand poses outperforms existing skeleton-based methods in Assembly101, incurring at least 5×5$\times$5 × fewer FLOPs. (iv) We experimentally demonstrate how hand poses are crucial for multiview and egocentric action recognition.
@@ 2 Related Work
Video Action recognition. Video-based action recognition systems are well-developed with sophisticated 3D-CNN [58, 9, 68, 23] or Video Transformer [2, 3, 48, 43] architectures. However, they all bear significant computational expense for both feature extraction and motion modeling, either explicitly in the form of optical flow [56, 9, 33] or implicitly through the architecture [2, 48]. Such designs are well-suited for high-facility applications but are not suitable for integration into lightweight systems. To this end, efficient video understanding is an active topic of research, with a focus on reducing expensive 3D operations [40, 68, 59, 22], quadratic complexity of attention [32, 48, 43] and dropping tokens [21, 4]. Additionally, densly-sampled RGB frames incur a high cost, yet limiting the temporal resolution will hinder fine-grained action understanding. We propose complementing 3D hand poses with sparsely-sampled RGB frames, developing a lightweight video understanding system.
@@
Skeleton-based Action Recognition. Skeleton-based action recognition has been tackled with hand-crafted features [63, 61], sequence models like RNNs and LSTMs [19, 70], and CNN-based methods that either employ temporal convolution on the pose sequence [57] or transform the skeleton data into pseudo-images to be processed with 2D or 3D convolutional networks [31, 7, 20]. Recent progress has been driven by GNN-based methods that exploit the graph structure of skeletal data to construct spatio-temporal graphs and perform graph convolutions [69, 71]. These methods often model functional links between joints that go beyond skeletal connectivity [39, 55] or expand the spatiotemporal receptive field [44]. Self-attention and transformer-based methods have also been proposed [49, 72, 66]. However, most existing methods are tailored for datasets involving full-body poses. Few works [30, 37, 51, 24] study hand poses but are restricted to simple gestures, recognizable without complex temporal modeling.
@@
Figure 2: Comparing skeletal changes in full-body actions from NTU RGB+D 120 [41] (left) and hand actions from Assembly101 [52] (right). Two pose frames at interval $T$ are shown. $J_{j,t}$ indicates the 3D coordinate of joint $j$ at timestep $t$, and ${<}J_{x},J_{y}{>}$ is the correlation between two such joints. Modeling the correlation between spatio-temporally distant joints can be informative for full-body poses but does not provide a useful action cue for hand poses.
@@
Fusing RGB with Skeleton. Skeletal data has been used for cropping images of body parts [12], for weighting RGB patches around regions of interest [5, 6], or for pooling CNN features [8, 1]. Projecting into common embedding space is done in [16], enabling pose distillation [15]. Multi-stream architectures are also designed with separate paths for RGB and skeleton having lateral connections between them [38, 64]. RGBPoseConv3D [20] is a two-stream model that employs 3D CNNs for both RGB and pose. Similar to skeleton-based action recognition, these multimodal approaches primarily focus on full-body poses. Some approaches simultaneously perform both hand pose estimation and action recognition, using pose data to supervise the training process [65, 13]. However, it is important to note that pose estimation deals with lower-level semantics than action recognition. Actions can often be inferred even when the estimated poses are not accurate [46].
3 Modeling Full-body vs. Hand Skeletons
@@
Existing skeleton-based action recognition datasets primarily consist of full-body poses where actions feature significant changes over time in limb positions relative to other body parts, which highly correlate to the action category. These changes can be captured through long-range spatiotemporal modeling using graph convolutions with large receptive fields [44] or self-attention [66]. Conversely, hand actions show diverse movement patterns, as hands can move in arbitrary directions, often without significant changes in articulation. We analyze pose sequences from NTU RGB+D 120 [41] and Assembly101 [52], calculating the distance covered by each joint and identifying the least and the most active joints. We observe that, for full-body poses, these two representative joints show a significant difference in the distances covered, while for hand poses, the difference is subtle. The Pearson correlation coefficient is 0.93 for hand poses (indicating highly coupled joints) and 0.33 for full-body poses. Please refer to Suppl. Sec. A for details regarding these computations and comparisons.
@@
An example showing the difference between body and hand poses is provided in Fig. 2, where we depict two frames separated by an interval of $T$ frames. For the full body skeleton, the head joint $J_{H}$ remains static, while the right hand joint $J_{F}$ moves upward. Conversely, for the hand, all the joints move, including the wrist joint $J_{W}$ and index fingertip $J_{I}$. The spatiotemporal correlation between the head joint in the first frame and the hand joint in the last frame (${<}J_{H,0},J_{F,T}{>}$) is a crucial action cue that captures the relative structural change. However, such a correlation between the wrist and the fingertip for hand skeleton (${<}J_{W,0},J_{I,T}{>}$) does not provide a stronger cue than the wrist movement itself. To avoid such redundant spatiotemporally distant correlations, we propose to divide the hand pose sequence into micro-action blocks of fixed temporal length. This formulation allows for encoding short-term movements while enabling parameter sharing.
@@
Moreover, full-skeletal motion is a dominant feature in hand pose sequences, differentiating them from full-body poses. While the location and orientation of the human body can be trivial for most actions, this is not the case for hands. The hands do not conform to any particular 6D pose, and notably, they frequently and unpredictably alter their 6D poses over time. A comprehensive analysis of this phenomenon is provided in Suppl. Sec. A. To this end, we explicitly consider the global 6D poses of the hands during micro-action-based pose encoding.
@@
Figure 3: Overall architecture of HandFormer. An action segment is divided into $K$ micro-actions $\{M_{1},M_{2}\ldots M_{K}\}$. Each micro-action comprises a dense sequence of pose frames and a single RGB frame. The frame encoder $F$ and the trajectory encoder $T$ encode the RGB and the dense poses, respectively, after which they are passed to a Multimodal Tokenizer. The modality-mixed tokens are then fed to a Temporal Transformer. Dotted paths are optional and only required when RGB is used.
@@ 4 HandFormer
Fig. 3 illustrates our proposed HandFormer, which consists of a sequence of micro-action blocks, a novel trajectory encoder, and a temporal aggregation module. The design of our HandFormer allows us to easily incorporate semantic context by sampling a single RGB frame from certain micro-actions.
@@ 4.1 Micro-actions
Given an action segment containing $\mathcal{T}$ frames sampled at a certain fps, the input to our model comprises — i) a dense sequence of 3D hand poses, represented as $\mathbf{S}=\{P_{1},P_{2},\ldots,P_{\mathcal{T}}\}$, where $P_{t}=\{P_{t}^{\text{{left}}},P_{t}^{\text{{right}}}\}$ $\in\mathbb{R}^{2\times J\times 3}$ signifying the 3D coordinates of $J$ keypoints in the left and right hands, respectively, and ii) a sparse set of RGB frames sampled at intervals $\Delta f$, denoted as $\mathbf{V}=\{I_{1},I_{1+\Delta f},\ldots,I_{\mathcal{T}_{r}}\}$, where $I_{t}\in\mathbb{R}^{H\times W\times 3}$ are frame-wise RGB features and $\mathcal{T}_{r}=1+\left\lfloor\frac{\mathcal{T}-1}{\Delta f}\right\rfloor\times\Delta f$.
@@ 4.2 Trajectory Encoder
We devise a trajectory-based pose encoder to encode dense hand pose sequences within micro-actions, illustrated in Fig. 4. Each joint is represented by its trajectory of dimension $3\times N$, encapsulating the sequence of 3D coordinates across the $N$ pose frames of a micro-action. This yields $2\times J$ feature vectors for the $J$ joints of two hands. Each joint’s trajectory is passed through a TCN [35], whose parameters are shared for all joints. This produces $2\times J$ Local Trajectory Tokens. Additionally, the full-skeletal motion of the hand during the entire action is used as a reference through an additional token named Global Wrist Token. This token is generated by a separate TCN operating on the sequence of 6D poses of hands, indicating wrist location and hand orientation. Subsequently, a self-attention layer is applied to these trajectory tokens, preserving the temporal dimension for subsequent stages. This iterative process culminates in spatiotemporal average pooling, summarizing the hand motion of the micro-action.
@@ 4.3 Multimodal Tokenizer
This section discusses incorporating sparsely sampled RGB frames into HandFormer to better capture scene semantics. A single frame is sampled from each micro-action, generating an extended crop (1.25×1.25$\times$1.25 ×) around the hands. While the full image provides an overall scene context, the crop focuses specifically on hand-object interaction regions [10]. Features for both are separately generated using a pre-existing image encoder and then aggregated to enrich the hand-object interaction feature with scene context. The hand-object ROI crop can be obtained using an off-the-shelf HOI detector [54] or by using the corresponding hand pose projections. We opt for the latter.
@@
Our multimodal tokenizer in Fig. 3 generates RGB and Pose tokens enhanced via multimodal interactions. This involves concatenating each frame feature and trajectory encoding, and projecting them into a shared PoseRGB feature space using an MLP. This PoseRGB feature is then split into two parts, which are added to the original frame feature and trajectory encoding, respectively.